OpenStudy (anonymous):

Create a difference of squares binomial or a perfect square trinomial that can be factored multiple times. (1 point) Part 2: Provide the factors of this polynomial. (2 points) Part 3: Explain, in complete sentences, the process you used to create the binomial/trinomial. (2 points)

OpenStudy (anonymous):

wow..um this is like something i wouldn't even try to do lol

OpenStudy (anonymous):

lolz i know right the teacher got me messed up

OpenStudy (anonymous):

lol i just passed the pretest.

OpenStudy (anonymous):

congrats :) I ddi too now im stuck on a test whos ur teacher

OpenStudy (anonymous):

(x + 1) × (x + 1) = x2 + x + x + 1 = x2 + 2x + 1 and x2 + 2x + 1 is a perfect square trinomial Another example is (x − 5) × (x − 5) (x − 5) × (x − 5) = x2 + -5x + -5x + 25 = x2 + -10x + 25 and x2 + -10x + 25 is a perfect square trinomial Now, we are ready to start factoring perfect square trinomials The model to remember when factoring perfect square trinomials is the following: a2 + 2ab + b2 = (a + b)2 and (a + b)2 is the factorization form for a2 + 2ab + b2 Notice that all you have to do is to use the base of the first term and the last term In the model just described, the first term is a2 and the base is a the last term is b2 and the base is b Put the bases inside parentheses with a plus between them (a + b) Raise everything to the second power (a + b)2 and you are done

OpenStudy (anonymous):

mrs.smith.....and your welcome :)

OpenStudy (anonymous):

@ilove mangoes Thx u very much U save my night :)

OpenStudy (anonymous):

@kiamonsta mr simpsom algebra 2 right honors or regular

OpenStudy (anonymous):

No problem. Just simple math.

OpenStudy (anonymous):

lol @iluvmangos im glad you find this simple!! lol

OpenStudy (anonymous):

2 ilove magoes that wasnt simple at all lolz

OpenStudy (anonymous):

@Samie509 i have honors

OpenStudy (anonymous):

Me too i have to finish the whole class by august 20 nd im only on chap 5 fml

OpenStudy (anonymous):

lol ohh snap. i have like 2 or 3 more mdules left

OpenStudy (anonymous):

Ur one lucky person @kiamonstaa

OpenStudy (anonymous):

lol not really. i have like three other classes to do also.

OpenStudy (anonymous):

ohh ok i got journalism 2 do so how long have ben in the class

OpenStudy (anonymous):

um..like since april

OpenStudy (anonymous):

Ohh u had time lolz i started in mid june

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

yeah